# When to use Graphs in RAG: A Comprehensive Analysis for Graph Retrieval-Augmented Generation

### v1

@@ Appendix D Evaluation Metrics
• ROUGE-L quantifies text similarity through n-gram overlap between generated and reference answers, capturing both syntactic and semantic alignment [40].
@@
{$FC&=2\cdot\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},\\ SS&=\text{cos}(\textbf{f}_{i},\textbf{c}_{j}),\end{aligned}\right.$ (6)
@@
Appendix E Additional Experiments
In our main experiments, we employ GPT-4o-mini as the default backbone model for generation. To evaluate how well different GraphRAG frameworks adapt across generation models, we use Qwen2.5-14B as the open-source generation component during generation. The experimental results are presented in Table 5. We summarize the observations as follows.
When integrated with the open-source Qwen2.5-14B model, several lightweight GraphRAG frameworks exhibit competitive performance. Notably, Fast-GraphRAG achieves the highest accuracy (60.08%) and ROUGE-L (41.31%) in fact retrieval, as well as strong performance in creative generation (ACC 47.60%) on the Novel dataset. On the Medical dataset, LightRAG leads in contextual summarization (ACC 69.37%) and creative generation (FS 70.84%), while HippoRAG2 obtains the best ROUGE-L scores for both fact retrieval (33.92%) and complex reasoning (33.02%). These results suggest that even under open-source settings, resource-efficient GraphRAG frameworks can effectively leverage graph-structured context to support various generation tasks.
Appendix F Implementation Details
F.1 Implementation Details of Representative RAG Models
@@
GraphRAG(local) [13] Textual Knowledge Graph Entity,Community Query Embedding Entity,Relationship,Chunk,Community Tabular Result
GraphRAG(global) [13] Textual Knowledge Graph Community Query Embedding Community(Layer) Literal Text
HippoRAG[20] Knowledge Graph Entity Entities in Query Chunk Reasoning path
HippoRAG2 [21] Knowledge Graph Phrase,Passage Query Embedding Phrase, Chunk Literal Text
LightRAG [19] Textual Knowledge Graph Entity,Relationship Keywords in Query Entity,Relationship,Chunk Literal Text + Graph Element
FastGraphRAG [11] Textual Knowledge Graph Entity Entities in Query Entity,Relationship,Chunk Literal Text
RAPTOR [48] Tree Treenode Query Embedding Tree node Reasoning path
As discussed in previous work  [16, 37, 65], RAG models comprise three core components: indexing, retrieval, and generation, each with its specific implementation details as presented in Table 8. Some explanation should be given to the content in the “Knowledge Type“ column of the table.A knowledge graph is constructed by extracting entities and relationships from each chunk, which contains only entities and relations, is commonly represented as triples. A textual knowledge graph is a specialized KG (following the same construction step as knowledge graph), which enriches entities with detailed descriptions and type information. A tree structure formed by document content and summary.
@@
  • GraphRAG(local): Microsoft-GraphRAG based on local retrieval granularity.
  • GraphRAG(global): Microsoft-GraphRAG based on global retrieval granularity.
@@
All of these GraphRAG methods construct graphs and refine retrieval strategies to boost RAG system performance across various specialized tasks. Due to time limits, we only assess several representative GraphRAG models. We will include more SOTA models, like ArchRAG [55] KET-RAG [30] PIKE-RAG [54] MedRAG [64] PathRAG [7] DBCopilot [56] LightPROF [2] CG-RAG [27].
F.2 The Specific Configuration of GraphRAG Models
@@
GraphRAG(global&local) Configuration
@@
Appendix G Related Work
G.1 Traditional RAG and their Limitations
The naive RAG systems [37] operate through three key steps: knowledge preparation, retrieval, and integration. During knowledge preparation, external sources such as documents, databases, or webpages are divided into manageable textual chunks and converted into vector representations for efficient indexing. In the retrieval stage, when a user submits a query, the system searches for relevant chunks using keyword matching or vector similarity measures. The integration stage then combines these retrieved chunks with the original query to create an informed prompt for the LLM’s response. Recent advancements in RAG systems have moved beyond basic text retrieval to structured knowledge integration. Modern implementations employ hierarchical architectures maintaining document organization via layered retrieval processes [8, 38], while others enhance precision through multi-phase retrieval mechanisms that first broaden then refine context selection [17, 59]. Autonomous query parsing frameworks automatically break down intricate questions into executable subqueries [4], complemented by context-aware systems that modify retrieval tactics in real-time according to query complexity and intent [52, 48]. These strategies advance naive RAG systems by improving context awareness, retrieval accuracy, and handling complex queries.
@@ G.1 Traditional RAG and their Limitations
Vector database architectures limit traditional RAG’s ability to handle multi-hop reasoning, as they retrieve information only from text chunks containing anchor entities. While methods like query expansion [43] and metadata enrichment [57] attempt to improve complex query handling, they remain constrained by the chunk-based knowledge structure that inherently disconnects related concepts. This structural limitation particularly hinders domain-specific reasoning requiring logical synthesis across distributed evidence. (ii) The chunking process sacrifices critical contextual relationships between specialized terms and abstract concepts, despite techniques like real-time retrieval alignment [33] and external API integration [36]. Vector databases’ flat organization fails to preserve hierarchical or conceptual dependencies essential for domain expertise utilization, leaving models unable to reconstruct professional knowledge networks from fragmented chunks. (iii) Vector similarity retrieval often overwhelms LLMs’ fixed context windows [44, 1] with redundant content, exacerbating their limited capacity for long-range dependency modeling. While strategies like context pruning [3] and LLM fine-tuning [41] reduce input volume, they cannot compensate for the structural inability to establish explicit connections between retrieved chunks. This fundamental mismatch persists despite optimizations like sliding windows [57], as vector-based approaches lack mechanisms for relational reasoning.
G.2 GraphRAG and its Advantages
@@
To address this, graph retrieval-augmented generation (GraphRAG) [45, 46] has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension [23, 61]. Early efforts, like Microsoft GraphRAG [13] and its variant LazyGraphRAG [12], employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG [19] improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG [28] introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG [38] tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG [39] constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
GraphRAG offers several significant advantages over traditional RAG systems [45], enhancing the capabilities of AI-powered information retrieval and generation. First, its graph-based knowledge representation captures hierarchical relationships and multi-hop dependencies between entities, enabling nuanced contextual reasoning and discovery of latent connections. This structured approach resolves ambiguity by evaluating multiple semantic paths during query processing. Besides, the graph structure allows unified integration of structured databases, semi-structured formats, and unstructured text within a single graph, supporting cross-modal queries that combine textual, numerical, and multimedia data. This interoperability maximizes value from heterogeneous organizational knowledge assets [46]. Third, users can audit decision pathways by visualizing entity relationships traversed during retrieval. Combined with LLMs, this transparent architecture supports multi-hop logical synthesis, critical for specialized domains like healthcare and finance [46, 23].
G.3 Existing Benchmarks and Analysis
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA [60], MultiHopRAG [53] and UltraDomain [47], fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. A few studies, like DIGIMON [65] and another analysis paper [22], have recently tried to analyze the effect of different GraphRAG models. Despite their effort, they mainly focus on architectural comparisons using homogeneous datasets, missing how models synthesize hierarchical expertise and unstructured narratives. To this end, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. It features hybrid corpora with tasks of increasing complexity and stage-specific metrics to expose why models fail, whether in graph construction, knowledge retrieval, or contextual synthesis. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
Appendix H Limitation
@@
Appendix I Broader Impact

### v2

@@ Appendix F Evaluation Metrics
ROUGE-L quantifies text similarity through n-gram overlap between generated and reference answers, capturing both syntactic and semantic alignment (Lin, 2004).
@@
$\left\{\begin{aligned} FC&=2\cdot\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},\\ SS&=\text{cos}(\textbf{f}_{i},\textbf{c}_{j}),\end{aligned}\right$. (6)
@@
Appendix G Additional Experiments
G.1 Experiments on more GraphRAG frameworks
We evaluated a total of $1$ GraphRAG frameworks in our study. Due to space constraints in the main paper, only a subset of the results is presented there. This appendix provides the complete and detailed results for all frameworks, which are shown in Table 5 and Table 6.
G.2 Experiments on open-source model
In our main experiments, we employ GPT-4o-mini as the default backbone model for generation. To evaluate how well different GraphRAG frameworks adapt across generation models, we use Qwen2.5-14B as the open-source model. The experimental results are presented in Table 7. Due to time and resource constraints, we tested a representative subset of these models on Qwen2.5-14B. We summarize the observations as follows.
When integrated with the open-source Qwen2.5-14B model, several lightweight GraphRAG frameworks exhibit competitive performance. Notably, Fast-GraphRAG achieves the highest accuracy (60.08%) and ROUGE-L (41.31%) in fact retrieval, as well as strong performance in creative generation (ACC 47.60%) on the Novel dataset. On the Medical dataset, LightRAG leads in Contextual Summarize (ACC 69.37%) and creative generation (FS 70.84%), while HippoRAG2 obtains the best ROUGE-L scores for both fact retrieval (33.92%) and complex reasoning (33.02%). These results suggest that even under open-source settings, resource-efficient GraphRAG frameworks can effectively leverage graph-structured context to support various generation tasks.
Appendix H Implementation Details
H.1 Implementation Details of Representative RAG Models
@@
MS-GraphRAG(local) (Edge et al., 2024) Textual Knowledge Graph Entity,Community Query Embedding Entity,Relationship,Chunk,Community Tabular Result
MS-GraphRAG(global) (Edge et al., 2024) Textual Knowledge Graph Community Query Embedding Community(Layer) Literal Text
HippoRAG (Gutiérrez et al., 2024) Knowledge Graph Entity Entities in Query Chunk Reasoning path
HippoRAG2 (Gutiérrez et al., 2025) Knowledge Graph Phrase,Passage Query Embedding Phrase, Chunk Literal Text
LightRAG (Guo et al., 2024) Textual Knowledge Graph Entity,Relationship Keywords in Query Entity,Relationship,Chunk Literal Text + Graph Element
FastGraphRAG (CircleMind-AI, 2024) Textual Knowledge Graph Entity Entities in Query Entity,Relationship,Chunk Literal Text
RAPTOR (Sarthi et al., 2024) Tree Treenode Query Embedding Tree node Reasoning path
KGP (Wang et al., 2024b) Knowledge Graph Entity, Relationship Query Graph Subgraph Linearized Subgraph
LAZY-GRAPHRAG Plain Text (on-demand graph) Text Chunk Query Embedding Chunk, Node (dynamic) Literal Text
structRAG (Li et al., 2024) Hierarchical Knowledge Graph Hierarchical Node Query Embedding Node, Path Literal Text + Structural Info
KET-RAG (Huang et al., 2025) Knowledge-Enhanced Tree Keyword, Entity, Chunk Query (multi-index) Keyword, Entity, Chunk Literal Text
As discussed in previous work  (Gao et al., 2023; Lewis et al., 2020; Zhou et al., 2025), RAG models comprise three core components: indexing, retrieval, and generation, each with its specific implementation details as presented in Table 10. Some explanation should be given to the content in the “Knowledge Type“ column of the table.A knowledge graph is constructed by extracting entities and relationships from each chunk, which contains only entities and relations, is commonly represented as triples. A textual knowledge graph is a specialized KG (following the same construction step as knowledge graph), which enriches entities with detailed descriptions and type information. A tree structure formed by document content and summary.
@@
  • MS-GraphRAG(local): Microsoft-GraphRAG based on local retrieval granularity.
  • MS-GraphRAG(global): Microsoft-GraphRAG based on global retrieval granularity.
@@
  • StructRAG: a framework that boosts knowledge-intensive reasoning of LLMs by dynamically restructuring scattered information into a hybrid, structured format at inference time.
  • KGP: a framework that improves multi-document question answering by constructing and traversing a knowledge graph to formulate the right context for large language models.
  • Lazy-GraphRAG: a framework that achieves a better cost-quality trade-off by using a "lazy" approach to build a concept graph only at query time.
  • KET-RAG: a framework that achieves a cost-efficient and high-quality Graph-RAG by leveraging a multi-granular indexing approach combining a knowledge graph skeleton with a text-keyword bipartite graph.
All of these GraphRAG methods construct graphs and refine retrieval strategies to boost RAG system performance across various specialized tasks. Due to time limits, we only assess several representative GraphRAG models. We will include more SOTA models, like ArchRAG (Wang et al., 2025b) PIKE-RAG (Wang et al., 2025a) MedRAG (Zhao et al., 2025) PathRAG (Chen et al., 2025) DBCopilot (Wang et al., 2025c) LightPROF (Ao et al., 2025) CG-RAG (Hu et al., 2025).
H.2 Configuration of GraphRAG Models
@@
  "chunk_token_size": 256,
MS-GraphRAG(global&local) Configuration
@@
Appendix I Related Work
I.1 Traditional RAG and their Limitations
The naive RAG systems (Lewis et al., 2020) operate through three key steps: knowledge preparation, retrieval, and integration. During knowledge preparation, external sources such as documents, databases, or webpages are divided into manageable textual chunks and converted into vector representations for efficient indexing. In the retrieval stage, when a user submits a query, the system searches for relevant chunks using keyword matching or vector similarity measures. The integration stage then combines these retrieved chunks with the original query to create an informed prompt for the LLM’s response. Recent advancements in RAG systems have moved beyond basic text retrieval to structured knowledge integration. Modern implementations employ hierarchical architectures maintaining document organization via layered retrieval processes (Chen et al., 2024a; Li et al., 2024), while others enhance precision through multi-phase retrieval mechanisms that first broaden then refine context selection (Glass et al., 2022; Xu et al., 2023). Autonomous query parsing frameworks automatically break down intricate questions into executable subqueries (Asai et al., 2023), complemented by context-aware systems that modify retrieval tactics in real-time according to query complexity and intent (Tang et al., 2024; Sarthi et al., 2024). These strategies advance naive RAG systems by improving context awareness, retrieval accuracy, and handling complex queries.
@@ I.1 Traditional RAG and their Limitations
Vector database architectures limit traditional RAG’s ability to handle multi-hop reasoning, as they retrieve information only from text chunks containing anchor entities. While methods like query expansion (Mao et al., 2020) and metadata enrichment (Wang et al., 2024a) attempt to improve complex query handling, they remain constrained by the chunk-based knowledge structure that inherently disconnects related concepts. This structural limitation particularly hinders domain-specific reasoning requiring logical synthesis across distributed evidence. (ii) The chunking process sacrifices critical contextual relationships between specialized terms and abstract concepts, despite techniques like real-time retrieval alignment (Jiang et al., 2024) and external API integration (Lazaridou et al., 2022). Vector databases’ flat organization fails to preserve hierarchical or conceptual dependencies essential for domain expertise utilization, leaving models unable to reconstruct professional knowledge networks from fragmented chunks. (iii) Vector similarity retrieval often overwhelms LLMs’ fixed context windows (OpenAI, 2023; Anthropic, 2024) with redundant content, exacerbating their limited capacity for long-range dependency modeling. While strategies like context pruning (Arefeen et al., 2024) and LLM fine-tuning (Luo et al., 2023) reduce input volume, they cannot compensate for the structural inability to establish explicit connections between retrieved chunks. This fundamental mismatch persists despite optimizations like sliding windows (Wang et al., 2024a), as vector-based approaches lack mechanisms for relational reasoning.
I.2 GraphRAG and its Advantages
@@
To address this, graph retrieval-augmented generation (GraphRAG) (Peng et al., 2024; Procko & Ochoa, 2024) has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension (Han et al., 2024; Zhang et al., 2025). Early efforts, like Microsoft GraphRAG (Edge et al., 2024) and its variant LazyGraphRAG (Darren Edge, 2024), employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG (Guo et al., 2024) improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG (Hu et al., 2024) introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG (Li et al., 2024) tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG (Liang et al., 2024) constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
GraphRAG offers several significant advantages over traditional RAG systems (Peng et al., 2024), enhancing the capabilities of AI-powered information retrieval and generation. First, its graph-based knowledge representation captures hierarchical relationships and multi-hop dependencies between entities, enabling nuanced contextual reasoning and discovery of latent connections. This structured approach resolves ambiguity by evaluating multiple semantic paths during query processing. Besides, the graph structure allows unified integration of structured databases, semi-structured formats, and unstructured text within a single graph, supporting cross-modal queries that combine textual, numerical, and multimedia data. This interoperability maximizes value from heterogeneous organizational knowledge assets (Procko & Ochoa, 2024). Third, users can audit decision pathways by visualizing entity relationships traversed during retrieval. Combined with LLMs, this transparent architecture supports multi-hop logical synthesis, critical for specialized domains like healthcare and finance (Procko & Ochoa, 2024; Han et al., 2024).
I.3 Existing Benchmarks and Analysis
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA (Yang et al., 2018), MultiHopRAG (Tang & Yang, 2024) and UltraDomain (Qian et al., 2024), fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. A few studies, like DIGIMON (Zhou et al., 2025) and another analysis paper (Han et al., 2025), have recently tried to analyze the effect of different GraphRAG models. Despite their effort, they mainly focus on architectural comparisons using homogeneous datasets, missing how models synthesize hierarchical expertise and unstructured narratives. To this end, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. It features hybrid corpora with tasks of increasing complexity and stage-specific metrics to expose why models fail, whether in graph construction, knowledge retrieval, or contextual synthesis. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
Appendix J Limitation
@@
Appendix K Broader Impact
@@
Appendix L The Use of Large Language Models (LLMs)
In the preparation of this manuscript, we used a large language model as a writing assistant. Its main role was to help improve our English writing, such as correcting grammar and refining sentences for clarity and style. Additionally, it was used to help set up the initial format for several tables. The authors made all final decisions on the content, carefully checking and editing all suggestions from the model to ensure the scientific accuracy and integrity of this work.